The perimeter of a rectangle is 96 inches. The length exceeds the width by 2 inches. Find the length and width.
Answer by checkley77(12569)

L=W+2
2W+2(W+2)=96
2W+2W+4=96
4W=96-4
4W=92
W=92/4
W=23IN. ANSWER FOR THE WIDTH.
L=23+2=25IN. ANSWER FOR THE LENGTH.
PROOF:
2*23+2*25=96
46+50=96
96=96